# Mixture of Routers

### v2

@@ 4.2 Main Results
Direct Fine-Tuning. We first compare the results of direct fine-tuning between MoR and baseline models on LLAMA-2-7B, where the accuracy (%) results of MoR and other baselines are shown in Table 1. We use the same hyperparameters for all methods. The results indicate that methods based on LoRA (LoRA, MoLA, and MoR) significantly outperform the baseline methods based on prompt tuning (Prompt Tuning and LLaMA-Adapter). After inserting the MoR module, there is a significant improvement across all tasks. Specifically, MoELoRA and MoLA show increased accuracy rates in the OpenbookQA task by 1.4% and 2.4%, respectively, and in the RTE task by 2.17% and 1.58%, respectively, after the integration of the MoR module. On average, the insertion of the MoR module has led to improvements of 1.03% and 0.98%, respectively. These results demonstrate that MoR can effectively enhance the performance of MoE-style models.
Figure 3: A visualization of expert allocation by MoR and MoLA, showing layers with a total of 4, 6, and 8 experts from left to right. Layers with only 2 experts are not displayed due to the use of the top-2 mechanism. The x-axis represents the expert IDs, and the y-axis represents the number of times each expert is activated.
Figure 4: Comparison of training loss between MoR and MoLA on the OpenBookQA task.
@@ 4.3.1 Router Number Analysis
Similar to MoE, MoR can also better adapt to different downstream tasks by adjusting the number of routers. In this section, we present the results and training times of MoR models with different numbers of routers across six distinct tasks, as shown in Table 2. The analysis shows that when using two routers ($r=2$) for joint allocation, the model achieves the fastest training speed and attains the best overall results, with an average improvement of 0.98% and only a 4.6% increase in training time. When using three routers ($r=3$) for joint allocation, the model achieves average suboptimal results; however, it is noteworthy that the model produces the best outcomes on the ScienceQA and MRPC tasks. Upon further increasing the number of routers to four ($r=4$), there is a decline in model performance, along with a 7.1% increase in training time. Further analysis indicates that ScienceQA and MRPC are the most complex tasks among similar types of three tasks, thus requiring more joint allocation experts through routing. In contrast, simpler tasks are prone to overfitting due to excessively high model complexity when using multiple routers.
@@ 4.3.2 Expert Allocation Analysis
In this section, we explain the reasons for the performance improvement after inserting the MoR plugin through expert allocation analysis. The changes in expert allocation after inserting MoR are shown in Figure 3. From the figure, we can observe that before inserting MoR, in the left plot, expert(1) is activated excessively frequently, while expert(3) is activated far less often, leading to an imbalanced model workload. Similar activation patterns can also be observed with other numbers of experts. However, after inserting MoR, the distribution of expert activations becomes much more balanced. The difference between the most frequently activated expert and the least frequently activated expert significantly decreases.
Regarding incorrect expert allocations, we can see from the loss curve in Figure 4 that the model with the MoR plugin shows a noticeable reduction in training loss. This indicates that multi-expert routing can better capture underlying patterns in the data, thereby improving the overall performance and generalization ability of the model. Specifically, the MoR plugin introduces a multi-routing mechanism, allowing the model to consider more information when selecting experts. This avoids over-reliance on a single expert or neglecting certain experts. This improvement in load balancing not only enhances training efficiency but also reduces potential biases caused by uneven expert allocation.
@@ 5 Conclusion
We propose a MoE-style plug-in named MoR, which effectively alleviates the issues of incorrect assignments and imbalances in expert allocation in MoE-style models through multi-routing joint allocation. This innovative, plug-and-play approach provides a flexible solution to the expert allocation problem in MoE models. Extensive experiments conducted on NLP and CR tasks show that MoR outperforms baseline models both in direct fine-tuning and instruction-based fine-tuning scenarios. As a plug-and-play PEFT, MoR can be applied to a wide range of tasks. Moreover, this work offers a promising research direction for enhancing MoE technology and PEFT methods.
Limitations
There are two limitations in this work. First, the current research primarily focuses on experiments with single-modality data and models. Our method has not yet been validated on models handling other modalities. Secondly, due to the lack of computational resources, the experiments in this study were limited to fine-tuning tasks and did not involve the process of training large-scale models from scratch. This means that the performance of our approach in training large models from the ground up has not been fully validated. We look forward to addressing these limitations in the future.
Appendix A Datasets
A.1 NLP Datasets
For evaluation, we adopt the GLUE (General Language Understanding Evaluation) benchmark, which is a widely used collection of datasets designed to assess natural language understanding (NLU) capabilities of models. Specifically, the GLUE benchmark includes several datasets such as COLA Wang et al. (2018), where the task is to determine whether a given sentence is grammatically acceptable; MRPC Dolan and Brockett (2005), which focuses on identifying whether two sentences are semantically equivalent; and RTE Wang et al. (2018), where the goal is to predict if a premise sentence entails a hypothesis sentence. We present the dataset statistics of GLUE in the following table 3.
Dataset Metric #Train #Valid #Test #Label
COLA Mcc 8.5k 1,043 1,063 2
MRPC Acc 3.7k 408 1.7k 2
RTE Acc 2.5k 277 3k 2
Table 3: Dataset Sizes and Evaluation Metrics in the GLUE Benchmark. "Mcc" and "Acc" denote the Matthews correlation coefficient and accuracy
A.2 CR Datasets
Specifically, for the CR tasks, we evaluate ScienceQA Lu et al. (2024), a dataset that tests multi-modal reasoning across text, images, and scientific concepts; CommonsenseQA Talmor et al. (2019), which focuses on answering questions that require everyday commonsense knowledge and reasoning about implicit relationships; and OpenbookQA Mihaylov et al. (2018), designed to evaluate reasoning over explicit scientific facts and principles provided in an open book format. For more details and specific statistics, please refer to Table 4.
Dataset Metric #Train #Valid #Test #Label
SciQA Acc 6,508 - 2,224 4
ComQA Acc 9,740 1,221 - 5
OpenQA Acc 4,957 500 500 4
Table 4: Dataset Sizes and Evaluation Metrics for Science Question Answering Tasks. "Acc" denotes accuracy.
Appendix B Redundancy and Fault Tolerance Theory
Redundancy and Fault Tolerance Theory is a fundamental concept in system design, particularly in engineering, computer science, and reliability analysis. It focuses on enhancing system robustness by incorporating redundant components or mechanisms to ensure continued operation despite failures. The theory is based on the principle that adding extra resources, such as backup systems or duplicate components, can mitigate the impact of faults, thereby improving overall reliability and availability. A key metric in this theory is the system reliability, often modeled using probability:
$R_{sys}=1-(1-R)^{n}$ (7)
where R represents the reliability of an individual component, and n is the number of redundant components. This formula illustrates how redundancy increases system reliability exponentially with additional components. Fault tolerance extends beyond hardware to software systems, employing techniques like error detection, correction codes, and failover mechanisms.
Appendix C Hyperparameters
To facilitate the reproducibility of our experimental results, we have made the hyperparameters used in our experiments publicly available. This includes detailed configurations such as learning rates, batch sizes, optimization algorithms, weight decay, and other relevant settings. By providing this information, we aim to ensure that our experiments can be replicated and validated by other researchers, fostering transparency and enabling further advancements in the field.
         Hyperparameters          Value
         batch_size          64
         micro_batch_size          4
         learning_rate          3e-4
         cutoff_len          256
         lora_r          8
         lora_dropout          0.05
         lora_alpha          16
Table 5: Hyperparameters
Figure 5: Visualization of MoR expert activations, with the total number of experts being 2, 4, 6, and 8 from top to bottom.
Appendix D Visualization of expert allocation
In this section, we present the visualization of expert allocation and activation within the MoR framework, as shown in Figure 5. The similarity in color across each row indicates that the MoR plugin effectively mitigates the issue of improper and uneven expert allocation commonly observed in MoE-style models.

### v3

@@ 4.2 Main results
4.2.1 Direct fine-tuning
We first compare the results of direct fine-tuning between MoR and baseline models on three commonly used models (Llama-7B-hf, Llama-2-7B-hf and Mistral-7B-v0.1). The accuracy results of MoR and other baseline models are shown in Table 1. We use the same hyperparameters for all methods. The results indicate that after inserting the MoR module, there is a significant improvement across most tasks for all models. Taking the Llama-2-7B-hf model as an example. After integrating the MoR module, MoELoRA and MoLA show increased accuracy rates in the OpenbookQA task by 1.4% and 2.4%, respectively, and in the RTE task by 2.17% and 1.58%, respectively. On average, the insertion of the MoR module has led to performance improvements of 1.03% and 0.98%, respectively. Consistent improvements are observed on the other two models, showing average performance gains of 0.71% and 0.44% after integrating MoR. These results demonstrate that MoR can effectively enhance the performance of MoE-style models.
4.2.2 Transfer fine-tuning
In the context of transfer fine-tuning, we first fine-tune LLAMA-2-7B-hf on an instructional tuning dataset using each PEFT (Parameter-Efficient Fine-Tuning) method. Subsequently, we further fine-tune the model on all downstream tasks. This instructional tuning process serves as an effective means to evaluate the transfer learning capabilities of each PEFT method. To ensure a focused comparison, we limit our analysis to methods based on LoRA, as these have demonstrated superior transfer learning performance compared to approaches reliant on prompt tuning. As shown in Table 2, MoR continues to outperform other methods across most benchmark tests. Notably, in the CommonsenseQA and OpenbookQA tasks, MoLA+MoR achieves significant improvements over MoLA, with performance gains of 2.9% and 0.82%, respectively. On average, the incorporation of MoR leads to a notable 1.16% improvement in overall performance. These experimental results underscore the value of MoR as a plug-and-play enhancement for MoE-style models. By integrating MoR, the overall performance of MoE models is significantly boosted, while also showcasing its robust transfer learning capabilities. This makes MoR a highly versatile and effective solution for a wide range of applications.
Table 2: The transfer learning capabilities of different models across various tasks. We use OpenOrca as the dataset for the first fine-tuning.
Models SciQA ComQA OpenQA MRPC CoLA RTE Avg.
LoRA 91.01 74.61 76.6 84.41 84.95 84.48 83.01
LoRAMoE 91.41 75.92 77.6 84.93 85.81 86.64 83.72
LoRAMoE + MoR (+0.95) (+1.23) (+0.80) (+0.00) (+0.48) (+1.09) 84.48 (+0.76)
MoLA 92.94 77.97 78.7 84.52 86.64 86.19 84.54
MoLA + MoR (+0.27) (+0.82) (+2.9) (+0.87) (+0.23) (+2.26) 85.70 (+1.16)
Table 3: The results of MoR with different router numbers were tested on 6 different tasks, and we report the accuracy and time efficiency. $r$ represents the number of routers, $h$ and $m$ represent hours and minutes, respectively.
Method Metric ComQA OpenQA CoLA RTE
MoLA Acc(%) 78.95 79.6 83.48 86.77
Time(train) 13.20h 9.67h 7.52h 15.80h
Time(test) 6.57m 2.53m 8.07m 1.50m
+ MoR Acc(%) 79.20 82.0 83.94 86.87
Time(train) 19.78h 10.03h 7.92h 17.00h
$(r=2)$ Time(test) 7.14m 2.55m 8.32m 1.53m
+ MoR Acc(%) 78.05 81.2 84.81 86.10
Time(train) 19.99h 10.27h 7.98h 17.15h
$(r=3)$ Time(test) 7.29m 3.05m 8.40m 1.55m
+ MoR Acc(%) 76.82 80.0 84.00 86.10
Time(train) 20.31h 10.57h 8.02h 17.31h
$(r=4)$ Time(test) 7.34m 3.11m 8.59m 2.05m
@@ 4.3.1 Router number analysis
Similar to MoE, MoR can also better adapt to different downstream tasks by adjusting the number of routers. In this section, we present the results and training times of MoR models with different numbers of routers across six distinct tasks, as shown in Table 3. The analysis shows that when using two routers ($r=2$) for joint allocation, the model achieves the best overall results, with an average improvement of 0.98%. When using three routers ($r=3$) for joint allocation, the model achieves average suboptimal results; however, it is noteworthy that the model produces the best outcomes on the ScienceQA and MRPC tasks. Upon further increasing the number of routers to four ($r=4$), there is a decline in model performance.
Further analysis indicates that ScienceQA and MRPC are the most complex tasks among similar types of three tasks, thus requiring more joint allocation experts through routing. In contrast, simpler tasks are prone to overfitting due to excessively high model complexity when using multiple routers. Moreover, unlike MoE, which requires dynamic selection among all experts for each input, MoR only needs to coordinate and make decisions among a small number of routers. Specifically, our results show that using a number of routers equivalent to approximately one-fourth the number of experts is sufficient to achieve near-optimal performance. This significantly reduces the complexity of routing decisions, as the main router in MoR only needs to assign weights among a few sub-routers. Consequently, MoR largely avoids the issue of imbalanced routing load that is commonly observed in MoE systems. In summary, the model performs best when using two routers ($r=2$) for joint allocation, achieving improved performance while maintaining high training efficiency.
4.3.2 Efficiency analysis
In this section, we conduct an analysis of the efficiency of MoR, with the relevant results presented in Table 3. The table provides a detailed breakdown of how different numbers of sub-routers impact training and inference efficiency. Experimental data shows that our best-performing model, MoLA+MoR $(r=2)$, introduces relatively low latency during both training and inference. Specifically, training time increases by an average of 4.41%, while inference time rises by an average of 2.58%. Although these additional costs are not negligible, they are entirely acceptable when compared to the significant performance improvements achieved by the model. Furthermore, the experimental results further validate the rationality of the model design, demonstrating that MoR can achieve a good balance between efficiency and performance when the number of sub-routers is moderate.
@@ 4.3.3 Expert allocation analysis
In this section, we evaluate how MoR affects expert allocation by recording the number of expert activations. The changes in expert allocation after inserting MoR are shown in Figure 2. Before inserting MoR, there was a significant difference between the most frequently activated experts (5, 7) and the least frequently activated experts (6, 8), reflecting an imbalance in workload distribution within the original MoE architecture. Similar activation patterns can also be observed under other configurations of the number of experts. However, after inserting MoR, the distribution of expert activations becomes more balanced, and the difference between the most and least activated experts is significantly reduced.
Figure 2: Visualizing the effect of incorporating the MoR plugin into the model through expert activation patterns with 8 experts.
Specifically, the MoR plugin introduces a multi-routing mechanism, allowing the model to consider more information when selecting experts. This avoids over-reliance on a single expert or neglecting certain experts. This improvement in load balancing not only enhances training efficiency but also reduces potential biases caused by uneven expert allocation.
@@ 5 Conclusion
In this work, we identify a critical yet underexplored limitation in current MoE based parameter-efficient fine-tuning (PEFT) approaches: the fragility of single-router mechanisms, which leads to misrouted expert assignments and severe load imbalance—especially as model scale and expert count grow. To address this, we propose Mixture of Routers (MoR), a novel plug-and-play PEFT framework that rethinks routing not as a monolithic decision, but as a collaborative, redundant process inspired by fault-tolerant system design.
MoR introduces multiple learnable sub-routers that jointly evaluate expert suitability, while a lightweight main router dynamically arbitrates their contributions. This multi-path routing strategy significantly enhances assignment accuracy and expert utilization balance, without incurring additional inference latency—a crucial requirement for real-world deployment. Extensive experiments across NLP and commonsense reasoning benchmarks demonstrate that MoR consistently outperforms strong MoE-LoRA baselines, achieving an average performance gain of 1% under both standard and instruction-based fine-tuning paradigms.
Acknowledgments
This work was supported in part by the Science and Technology Commission of Shanghai Municipality under Grant (21DZ2203100), in part by the National Natural Science Foundation of China under Grant (62006150).
Data availability
The data used in this study are publicly available.
[…]